INAF LBC+LUCIFER queue observing, 2010 October 09-10

Observer: V. Testa, K.Boutsia, M. Fumana
Telescope Operator: D.Gonzales Huerta
Support Astronomer: O.Kuhn
Instrument: LBC - LUCIFER

Summary

relatively clear sky with some clouds during the night
Seeing on average decent: ~1" in the IR, ~1.2-1.5 in R (according to the AG FWHM)
30min lost because of telescope communication problems
30min lost because of instrument problems (LBC)

Data Summary

Project Objects FilterSorted ascending ExpTime Notes
Longhetti S2F1-142 150Ks 4hr seeing below 1" -Telluric+ARC+FLAT, maybe some clouds
Turatto SN2009jf J, H, Ks 80s + 160s + 240s seeing around 1"
Cresci DSF2237A Ks 50min seeing below 1", maybe 5min with clouds

Details

01:17 - Strange flats

01:48 - Slew to pointing star ACT0219
  • Pointing correction OK + mirror adjust OK

01:51 - slew to SA113 dofpia,
  • dofpia,/First
  • red image looks like a bias.

02:03 - Turn OFF/ Turn ON LBC
  • Does not work. It's a problem with the r-SLOAN filter wheel, try other filters

02:35 - Use a last-but-one version of the software
  • No light from r-SLOAN as well. Problem with filter wheels in the red.
  • Seeing is however quite poor (~1.5")

02:40 - SWITCH TO LUCIFER

03:11 - Slew to SN2009jf image acquisition (Turatto) 001-002

03:18 - Science imaging scripts for SN2009jf
  • Imaging H 003-004
  • Imaging J 005-008
  • OB interrupted because objects are elongated - INITIALIZE

03:29 - Image acquisition for SN2009jf 011-012
  • changed acquisition script offset to absolute
  • Science imaging script H 013-016
  • Science imaging script J 017-020
  • Science imaging script Ks 021-024

03:47 - Slew to CampoA (Cresci)
  • NO GUIDE STAR

03:55 - Slew to CampoA (Cresci) after changing guide star 025-028
  • interrupted because images are elongated again
  • To correct focus, we switch cameras (N3.75 -> N1.8 -> N3.75) and INITIALIZE
  • restart OB - interrupted because of ReadOut Manager error - IntegrationAbortedException

04:10 - Science script for Ks Imaging CampoA 0029-0073
  • 45 Jitters done, exited at 05:27 by telescope comm. error

05:32 - Connected again the telescope but still not working, restart telescope server on the Lucifer side

05:39 - Back again to Cresci Campo A collimating

05:50 - CLOUDS!!
  • Lost target, script stopped. 0074-0081 probably last two or three not good

05:56 - Slew to Telluric1 Longhetti (error in committing telescope)
  • Unable to get DD entries via IRS (error message)

06:28 - Telescope problem solved - slew to TEL1 (Longhetti)
  • acquisition 0082-0084

06:36 - Science script fot TEL1 0085-0092

06:41 - Slew to science target Longhetti
  • acquisition 0093-0095

06:54 - Science script Longhetti 8th run 0096-0103
  • at the end of the OB offset Y +64 to go back to starting position

07:38 - Science script Longhetti 9th run 0104-0111
  • at the end of the OB offset Y +64 to go back to starting position

08:23 - Science script Longhetti 10th run 0112-0119
  • at the end of the OB offset Y +64 to go back to starting position

09:07 - Science script Longhetti 11th run 0120-0127
  • at the end of the OB offset Y +64 to go back to starting position

09:52 - Science script Longhetti 12th run 0128-0135
  • at the end of the OB offset Y +64 to go back to starting position

10:36 - Science script Longhetti 13th run 0136
  • on 2nd offset - lost guiding - OB canceled

10:47 - repeat acquisition of Longhetti target 0137-0139

11:02 - Repeat science script Longhetti 13th run 0140-147

11:47 - Slew to Telluric2 Longhetti
  • target acquisition 148-152

12:00 - Science script fot TEL2 0153-0160

12:04 - Slew back to science target Longhetti because images of the last run seem faint
  • clouds? - out of focus? - target not centered in the slit ?
  • repeat acquisition and last run after INITIALIZE
  • target acquisition 0161-0163

12:21 - Repeat science script Longhetti 13th run 0164-167
  • after 4th integration OB interrupted because of twilight

12:44 - Calibration Unit IN
  • Start with FLATS - aborted 0168-0192
  • ARCS with flexure compensation OFF 0193-0198
  • ARCS with flexure compensation ON 0199-0204
  • These calibration files DO NOT have PARTNER and PI on the header
  • correct script - including archive information
  • move telescope to the HORIZON
  • FLATS with flexure compensation ON and telescope at the horizon 0205-0454

13:35 - Darks - End of Night

-- VincenzoTesta - 10 Oct 2010
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback